OpenStudy (anonymous):

8h-10h=3h+25 solve for h?

OpenStudy (anonymous):

-2h-3h=25 -5h=25 h=-5

OpenStudy (anonymous):

8h - 3h -10h = 25 => -5h = 25 => h = 25/(-5) => h = -5

OpenStudy (anonymous):

Thank you. But how did you get that answer?

OpenStudy (anonymous):

u didnt get the solution? mean the simplifications?

OpenStudy (anonymous):

come to www.chatindia.co.in and i'll tell you there. due to this slow site i can't help you fast.

OpenStudy (anonymous):

I can't leave this website- Chatindia and I just dont knw how to solve it. I got the answer. - uzma

OpenStudy (anonymous):

u need some explainations?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

look...on the right side we have 8h-10h what does it give?

OpenStudy (anonymous):

oh on the left :)

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

-2h,right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now we have got one term involving h on the right side shift it to left,u would have change in sign

OpenStudy (anonymous):

so it comes -2h-3h=25 right?

OpenStudy (anonymous):

what about -2h=3h+25

OpenStudy (anonymous):

shift 3h to the left

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then -2h-3h=25 gives -5h=25 ok?

OpenStudy (anonymous):

ok thanks :D

OpenStudy (anonymous):

so dividing both sides by -5,we get h=-5 :)

OpenStudy (anonymous):

so i have another. i just want to get this stuff down. X)

OpenStudy (anonymous):

y+2(y-5)=2y+2

OpenStudy (anonymous):

oh sure

OpenStudy (anonymous):

y +2y-10=2y+2 got it?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

y+2y-2y=2+10 y=12

OpenStudy (anonymous):

again collecting the term involving y on one side,and constant on the other

OpenStudy (anonymous):

should i subtract 2y from the other side?

OpenStudy (anonymous):

from the right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

well it would be cancelled out with +2y

OpenStudy (anonymous):

aren you supposed to cancle everything out?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok do one thing u try a solution,i wanna check where r u facing the prob:)

OpenStudy (anonymous):

y=2+10y=12 ?I dont know :o

OpenStudy (anonymous):

no solve the left side first n post it

OpenStudy (anonymous):

start from the begining?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

y^2-3 is the left side i got

OpenStudy (anonymous):

y+2(y-5) is the left side ,ok?

OpenStudy (anonymous):

2 is multiplied with both the terms in parenthesis

OpenStudy (anonymous):

n we get y+2y-10 right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now first simplify it,n post

OpenStudy (anonymous):

the left?

OpenStudy (anonymous):

3y-10

OpenStudy (anonymous):

fine :)

OpenStudy (anonymous):

now whats on the right involving y? shift it to the left, n post

OpenStudy (anonymous):

3y-10+2y=2

OpenStudy (anonymous):

good show, simplify the term involving y n shift 10 on the right

OpenStudy (anonymous):

o o, when u shift 2y on the left, it becomes -2y

OpenStudy (anonymous):

5y=-8?

OpenStudy (anonymous):

nops

OpenStudy (anonymous):

read my previous

OpenStudy (anonymous):

it must be 3y-2y+10=2,right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

ah wrong me ...sorry 3y-2y-10=2

OpenStudy (anonymous):

that makes more sence!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

now simplify it

OpenStudy (anonymous):

1y-10=2

OpenStudy (anonymous):

shift 10 to the right

OpenStudy (anonymous):

u get y=12 :)

OpenStudy (anonymous):

iy=12

OpenStudy (anonymous):

1y=12

OpenStudy (anonymous):

1y=y?

OpenStudy (anonymous):

THEN Y=12 OH CAPS..

OpenStudy (anonymous):

u got it now?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

fine :)any more prob?

OpenStudy (anonymous):

no(:

OpenStudy (anonymous):

y :(,cheer up :)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

nothing, take care:)

OpenStudy (anonymous):

(: